# Cyclic Counterfactuals under Shift–Scale Interventions

### v1

@@ 2.2 Semantics of Cyclic SCMs
One natural interpretation of cyclic structural equations is by assuming an underlying discrete-time dynamical system, where the equations act as update rules: the value of each variable at time t+1t{+}1 is computed from the values at time $t$. The system is then analyzed in the limit as $t\to\infty$, focusing on the fixed points to which the dynamics converge. Mooij et al. (2013) demonstrate that an alternative, yet natural, interpretation of SCMs emerges when considering systems of ordinary differential equations (ODEs). By examining the equilibrium (steady-state) solutions of such ODEs, one arrives at a structural causal model that is time-independent, but still retains meaningful causal semantics with respect to interventions. Specifically, the semantics of interventions and counterfactuals (see also Appendix A.1) remain valid and well-defined in this steady-state context, as rigorously formalized by Mooij et al. (2013) and further extended by Bongers et al. (2021). These are by no means the only routes to structural causal models; indeed, SCMs may arise through a variety of alternative constructions and representations, depending on the nature of the system under consideration. Although many physical processes exhibit inertia, static cyclic structural causal models (SCMs) remain appropriate when we focus on equilibrium behavior or sample at a temporal resolution coarser than the fastest feedback loop. Examples include gene‑regulatory networks in single‑cell genomics (Rohbeck et al., 2024); market‑equilibrium models (Bongers et al., 2021); predator–prey ecological systems; Thyroid or reproductive hormone axes exhibit feedback loops (Clarke et al., 2014), etc. Our mathematical framework is agnostic about the interpretation of cycles: we formulate everything directly at the level of structural equations with exogenous noise.
@@ 3 Theory
where $\mathbf{E}\sim\mathbb{P}_{\mathcal{E}}$ is the exogenous random variable. By the global $\kappa$-contraction in $x$ and the $1$-Lipschitz property in $e$ (both in $\ell^{2}$), $\Phi$ is $L$-Lipschitz with constant $L:=\frac{1}{1-\kappa}$ in $\ell^{2}$:
$\|\Phi(e_{1})-\Phi(e_{2})\|_{2}\leq\frac{1}{1-\kappa}\|e_{1}-e_{2}\|_{2}\quad\text{for all }e_{1},e_{2}\in\mathcal{E}.$
@@
$\mathbb{P}\Big(h(\mathbf{X},\mathbf{X}^{\prime})-\mathbb{E}[h(\mathbf{X},\mathbf{X}^{\prime})]\geq t\Big)\leq\exp\!\left(-\frac{t^{2}}{2L^{2}\sigma^{2}}\right)=\exp\!\left(-\frac{t^{2}}{2(1-\kappa)^{-2}\sigma^{2}}\right).$
@@
$\mathbb{P}\!\Big(h(\mathbf{X},\mathbf{X}^{\prime})-\mathbb{E}h(\mathbf{X},\mathbf{X}^{\prime})\geq t\Big)\;\leq\;\exp\!\Big(-\tfrac{t^{2}}{2\,(1-\kappa)^{-2}\,\|\Sigma^{1/2}\|_{2\to p}^{\,2}}\Big),$
@@
so the proxy equals $(1-\kappa)^{-2}\sigma^{2}$ for $p\geq 2$, and $(1-\kappa)^{-2}\sigma^{2}\,d^{\,2(\frac{1}{p}-\frac{1}{2})}$ for $1\leq p<2$. Proof sketch. Write $\mathbf{E}=\mu+\Sigma^{1/2}\mathbf{Z}$, $\mathbf{Z}\sim\mathcal{N}(0,I)$; then $z\mapsto h(\Phi(\mu+\Sigma^{1/2}z))$ is $(1-\kappa)^{-1}\|\Sigma^{1/2}\|_{2\to p}$-Lipschitz in $\ell^{2}$, and Gaussian Lipschitz concentration applies.
@@ 6 Conclusion and Future Work
Appendix A Supplementary Material
A.1 Equivalence of Twin SCM and Pearl’s Abduction-Action-Prediction (AAP) Procedure
Theorem 3.
Let $\mathcal{M}=\langle\mathcal{I},\mathcal{J},\mathcal{X},\mathcal{E},f,\mathbb{P}_{\mathcal{E}}\rangle$ be a structural causal model that is simple (i.e., uniquely solvable with respect to every subset of endogenous variables). Assume: (i) the exogenous vector $\mathbf{E}\sim\mathbb{P}_{\mathcal{E}}$ has a joint density and mutually independent components; (ii) the counterfactual query is defined by a perfect intervention $\mathrm{do}(X_{\mathcal{A}}:=\tilde{x}_{\mathcal{A}})$ for some $\mathcal{A}\subseteq\mathcal{I}$. Let $\mathbf{X}\sim\mathcal{M}$ denote the factual world, and define the counterfactual outcome via: Twin SCM definition: Solve the intervened twin SCM where the do-intervention is applied to the primed copy, yielding random variables $(\mathbf{X},\mathbf{X}^{\prime})$. Pearl’s AAP (Pearl_2009) procedure: Sample $\mathbf{E}\sim\mathbb{P}_{\mathcal{E}}$ conditioned on $\mathbf{X}=x^{\mathrm{obs}}$, and solve the intervened SCM with this fixed noise to obtain $\mathbf{X}^{\text{cf}}$. Then, the counterfactual distribution derived from the twin SCM coincides with the distribution obtained from Pearl’s abduction–action–prediction (AAP) procedure: $\mathbb{P}_{\mathbf{X}^{\prime}\mid\mathbf{X}=x^{\mathrm{obs}}}\;=\;\mathbb{P}_{\mathbf{X}^{\mathrm{cf}}\mid\mathbf{X}=x^{\mathrm{obs}}}.$
Proof.
Since $\mathcal{M}$ is simple, by definition there exists a unique measurable function $g:\mathcal{E}\to\mathcal{X}$ such that for every $e\in\mathcal{E}$,
[…]
Similarly, after the intervention $\mathrm{do}(X_{\mathcal{A}}:=\tilde{x}_{\mathcal{A}})$, the intervened SCM admits a unique measurable solution map $g^{\mathrm{do}}:\mathcal{E}\to\mathcal{X}$ (Proposition 3.8, Bongers et al. (2021)).
In the twin SCM construction, both copies (factual and counterfactual) share the same exogenous vector $\mathbf{E}$, and the intervention is applied only to the first copy. Solving the twin SCM yields:
$(\mathbf{X},\mathbf{X}^{\prime})=(g(\mathbf{E}),\;g^{\mathrm{do}}(\mathbf{E})).$
Let $x^{\mathrm{obs}}\in\mathcal{X}$ be the observed factual outcome. Pearl’s procedure involves:
  • Abduction: condition on the observation $\mathbf{X}=x^{\mathrm{obs}}$, which corresponds to conditioning on the set $\{e\in\mathcal{E}:g(e)=x^{\mathrm{obs}}\}$;
  • Action: apply the intervention to obtain the new function $g^{\mathrm{do}}$;
  • Prediction: evaluate $\mathbf{X}^{\mathrm{cf}}=g^{\mathrm{do}}(\mathbf{E})$ using the same noise, but now drawn from the posterior $\mathbb{P}_{\mathcal{E}\mid g(\mathbf{E})=x^{\mathrm{obs}}}$.
Define the posterior distribution on exogenous noise:
$\mu_{x^{\mathrm{obs}}}(A)\;:=\;\mathbb{P}\bigl(\mathbf{E}\in A\,\big|\,g(\mathbf{E})=x^{\mathrm{obs}}\bigr),\qquad\text{for all measurable }A\subseteq\mathcal{E}.$
This measure is supported on the set
$\{e\in\mathcal{E}:g(e)=x^{\mathrm{obs}}\}.$
Equality of counterfactual laws.
By construction, $\mathbf{X}^{\prime}=g^{\mathrm{do}}(\mathbf{E}),$ so the counterfactual distribution from the twin SCM (given $\mathbf{X}=x^{\mathrm{obs}}$) is:
$\mathbb{P}_{\mathbf{X}^{\prime}\mid\mathbf{X}=x^{\mathrm{obs}}}(B)\;=\;\mu_{x^{\mathrm{obs}}}\bigl(e:g^{\mathrm{do}}(e)\in B\bigr),\qquad B\subseteq\mathcal{X}.$
Pearl’s method also conditions on $\{\mathbf{X}=x^{\mathrm{obs}}\}$, inducing the same posterior $\mu_{x^{\mathrm{obs}}}$ over exogenous variables. Then, the counterfactual outcome is $\mathbf{X}^{\mathrm{cf}}=g^{\mathrm{do}}(\mathbf{E})$ with $\mathbf{E}\sim\mu_{x^{\mathrm{obs}}}$. Thus,
$\mathbb{P}_{\mathbf{X}^{\mathrm{cf}}\mid\mathbf{X}=x^{\mathrm{obs}}}(B)\;=\;\mu_{x^{\mathrm{obs}}}\bigl(e:g^{\mathrm{do}}(e)\in B\bigr),$
which is the same expression as for the twin SCM. ∎ The counterfactual law derived from the twin SCM coincides with the one obtained from Pearl’s abduction–action–prediction procedure under unique solvability.
Remark 3 (Notation $\mathbf{X}\sim\mathcal{M}$).
The shorthand $\mathbf{X}\sim\mathcal{M}$ means that the random vector $\mathbf{X}$ is a solution of the SCM $\mathcal{M}$.

### v2

@@ 2.2 Semantics of Cyclic SCMs
One natural interpretation of cyclic structural equations is by assuming an underlying discrete-time dynamical system, where the equations act as update rules: the value of each variable at time t+1t{+}1 is computed from the values at time $t$. The system is then analyzed in the limit as $t\to\infty$, focusing on the fixed points to which the dynamics converge. Mooij et al. (2013) demonstrate that an alternative, yet natural, interpretation of SCMs emerges when considering systems of ordinary differential equations (ODEs). By examining the equilibrium (steady-state) solutions of such ODEs, one arrives at a structural causal model that is time-independent, but still retains meaningful causal semantics with respect to interventions. Specifically, the semantics of interventions and counterfactuals (see also Appendix 7.1) remain valid and well-defined in this steady-state context, as rigorously formalized by Mooij et al. (2013) and further extended by Bongers et al. (2021). These are by no means the only routes to structural causal models; indeed, SCMs may arise through a variety of alternative constructions and representations, depending on the nature of the system under consideration. Although many physical processes exhibit inertia, static cyclic structural causal models (SCMs) remain appropriate when we focus on equilibrium behavior or sample at a temporal resolution coarser than the fastest feedback loop. Examples include gene‑regulatory networks in single‑cell genomics (Rohbeck et al., 2024); market‑equilibrium models (Bongers et al., 2021); predator–prey ecological systems; Thyroid or reproductive hormone axes exhibit feedback loops (Clarke et al., 2014), etc. Our mathematical framework is agnostic about the interpretation of cycles: we formulate everything directly at the level of structural equations with exogenous noise.
@@ 3 Theory
where $\mathbf{E}\sim\mathbb{P}_{\mathcal{E}}$ is the exogenous random variable. By the global $\kappa$-contraction in $x$ and the $1$-Lipschitz property in $e$ (both in $\ell^{2}$), $\Phi$ is $L$-Lipschitz with constant $L:=\frac{\sqrt{2}}{1-\kappa}$ in $\ell^{2}$:
$\|\Phi(e_{1})-\Phi(e_{2})\|_{2}\leq\frac{\sqrt{2}}{1-\kappa}\|e_{1}-e_{2}\|_{2}\quad\text{for all }e_{1},e_{2}\in\mathcal{E}.$
@@
$\mathbb{P}\Big(h(\mathbf{X},\mathbf{X}^{\prime})-\mathbb{E}[h(\mathbf{X},\mathbf{X}^{\prime})]\geq t\Big)\leq\exp\!\left(-\frac{t^{2}}{2L^{2}\sigma^{2}}\right)=\exp\!\left(-\frac{t^{2}}{4(1-\kappa)^{-2}\sigma^{2}}\right).$
@@
$\mathbb{P}\!\Big(h(\mathbf{X},\mathbf{X}^{\prime})-\mathbb{E}h(\mathbf{X},\mathbf{X}^{\prime})\geq t\Big)\;\leq\;\exp\!\Big(-\tfrac{t^{2}}{2^{1+\frac{2}{p}{}}\,(1-\kappa)^{-2}\,\|\Sigma^{1/2}\|_{2\to p}^{\,2}}\Big),$
@@
so the proxy equals $2^{\frac{2}{p}}(1-\kappa)^{-2}\sigma^{2}$ for $p\geq 2$, and $2^{\frac{2}{p}}(1-\kappa)^{-2}\sigma^{2}\,d^{\,2(\frac{1}{p}-\frac{1}{2})}$ for $1\leq p<2$. Proof sketch. Write $\mathbf{E}=\mu+\Sigma^{1/2}\mathbf{Z}$, $\mathbf{Z}\sim\mathcal{N}(0,I)$; then $z\mapsto h(\Phi(\mu+\Sigma^{1/2}z))$ is $2^{\frac{1}{p}}(1-\kappa)^{-1}\|\Sigma^{1/2}\|_{2\to p}$-Lipschitz in $\ell^{2}$, and Gaussian Lipschitz concentration applies.